OpenStudy (anonymous):

How can an expression written in either radical form or rational exponent form be rewritten to fit the other form?

OpenStudy (anonymous):

roots and exponents are "inverses" of sorts. for example \(\large \sqrt[4]{x^4}=x\\\huge \sqrt[\cancel{4}]{x^\cancel{4}}=x\)

OpenStudy (anonymous):

also recall the property \(\huge \sqrt[n]{x}=x^{\frac{1}{n}}\)